I re-aligned the beam onto the MC TRANS QPD since Kiwamu had centered the spots on the mirrors. However, I then inspected the MC2F camera. After coming back into the control room I noticed that the MC transmission had gone down by 50% and that the MC2 OSEMs showed a large step. My guess is that somehow the opening and closing of the can shifted the suspension. So I adjusted the MC2 alignment biases to recover the transmitted power (its now ~50000 instead of the ~33000 from Friday).